OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

Distribute first then combine like terms.

hero (hero):

Or you could do this: 3(x + 1) = -2(x - 1 + 2) 3(x + 1) = -2(x + 1) 3(x + 1) + 2(x + 1) = 0 5(x + 1) = 0 (x + 1) = 0/5 x + 1 = 0 x = ?

OpenStudy (anonymous):

what @Hero said haha.. also what @The_Muslim_Ninja said works too :P whenever you get problems asking you to solve for a variable you first need to simplify what you can. and as long as you're working towards isolating that variable, then you should be fine with either method :)

OpenStudy (nirmalnema):

1st open the brackets on both the sides then u will get 3x+3=-2x+2-4 then transfer the like terms to the same side so u get 3x+2x=2-4-3 5x=-5 so x =-1